OpenStudy (anonymous):

If g(x) = 2x- x + 5, then g (2)=

OpenStudy (anonymous):

@Callisto can you help me plz

OpenStudy (callisto):

Oh... it's better for you to close the question and re-post the question in the right group - mathematics group

OpenStudy (callisto):

And for your question g(x) = 2x- x + 5 = x+5 g(2) = (2) +5 = 7 My bad :|

OpenStudy (anonymous):

g(2) is read as "g of 2" and means to plug 2 into the g(x) equation wherever you see an x.

OpenStudy (anonymous):

so it 7 and wow i did not know this wasnt maths

OpenStudy (anonymous):

oh ok thanks smoothmath